OpenStudy (anonymous):

What degree Fahrenheit is it if the temperature is 15 degrees Celsius? can someone show all steps or reasoning thank you

OpenStudy (karatechopper):

15-32=-17 -17 x 5=-85 -85/9=-9.44

OpenStudy (karatechopper):

hope thts right

OpenStudy (turingtest):

You have been given the formula I assume?

OpenStudy (anonymous):

First you need to know the relationship between Fahrenheit and Celsius F = C x (9/5) + 32 Since we have 15 degrees Celsius, the Fahrenheit is given by F = 15 x (9/5) + 32 = 59 Thus, we have 59 degrees Fahrenheit

OpenStudy (karatechopper):

or its 15 x 1.8=27 27+ 32=59 tht seem s better^

OpenStudy (anonymous):

thank you so much

OpenStudy (karatechopper):

welcs please be sure to click good answer for us all

OpenStudy (anonymous):

I will promise thanks again

OpenStudy (karatechopper):

np